Calculate the volume (in dm3) of 1.50 mol.dm-3 H2SO4 necessary to react completely with 2.000 g CuO according to the equation below.

 

CuO + H2SO4 → CuSO4 + H2O


1
Expert's answer
2021-04-13T02:07:59-0400

n = m/M

M (CuO) = 79.545 g/mol

n (CuO) = 2.000 / 79.545 = 0.025 mol

C = n / V

V = n / C

According to the equation, n (CuO) = n (H2SO4).

V (H2SO4) = 0.025 / 1.5 = 0.017 L or dm3